The shape of a barbeque pit is like a parallelogram with height of 22 in and a base of 48 in. It costs $1.20 per ft^2 to fill the pit with charcoals. How much totally does it cost?

Answer by Alan3354(21594)

The shape of a barbeque pit is like a parallelogram with height of 22 in and a base of 48 in. It costs $1.20 per ft^2 to fill the pit with charcoals. How much totally does it cost?
------------------
Area = 22*48 = 1056 sq inches = (1056/144) sq ft
$1.20*1056/144 = $8.80